OpenStudy (anonymous):

Rectangles A and B are similar rectangles. The length of the diagonal of Rectangle A is 13 inches, and the length of the diagonal of Rectangle B is 6.5 inches. What could be the length and width of both Rectangle A and Rectangle B? Rectangle A: 5 in x 12 in, Rectangle B: 2.5 in x 6 in Rectangle A: 4 in x 10 in, Rectangle B: 3 in x 7 in Rectangle A: 7 in x 11 in, Rectangle B: 2 in x 5 in Rectangle A: 6.5 in x 14 in, Rectangle B: 3.5 in x 8 in

OpenStudy (anonymous):

\[13^2=5^2+12^2\]because\[169=25+144\] The ratio of the sides are 13:6.5=2:1 \[5*1/2=2.5 \]\[12*1/2=6\] 1st one.

OpenStudy (anonymous):

Thank you ! :)